tayjah:

Mr. Harding uses 5/2 cans of beans to make 3 servings of chili for dinner. Find the unit rate to figure out how many cans of beans he would need to make 1 serving of chili.

jhonyy9:

5/2 to 3 servingce x to 1 serv x = 5,2*1/3 = ?

BrotherBear:

5/2 divided by 3. this should get you 0.833

tayjah:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 5/2 to 3 servingce x to 1 serv x = 5,2*1/3 = ? \(\color{#0cbb34}{\text{End of Quote}}\) Huh

tayjah:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BrotherBear 5/2 divided by 3. this should get you 0.833 \(\color{#0cbb34}{\text{End of Quote}}\) Ok thx

BrotherBear:

Yup and if that has to be a fraction it would be 83/100

tayjah:

Ok

tayjah:

THX!😁